OpenStudy (anonymous):

The surface area of a rectangular prism is 42 cm2. What is the surface area of this prism, measured in mm2? A. 420 mm2 B. 42,000 mm2 C. 0.42 mm2 D. 4200 mm2

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not C.

OpenStudy (skullpatrol):

$$\huge cm^2 =(cm)^2$$

OpenStudy (anonymous):

surface area of a rectangular prism is 42 cm^2 recall 1 cm = 10 mm convert the units. 42 cm^2 = 42 * (10 mm)^2 = 42 * 100 mm^2 =4200 mm^2 i think

OpenStudy (anonymous):

i'm not that smart

OpenStudy (skullpatrol):

Correct.

OpenStudy (anonymous):

thank you

OpenStudy (skullpatrol):

Thanks for trying to understand :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

The volume of a prism is 2600 mm3. What is the volume of this prism, measured in cm3? A. 2.6 cm3 B. 26,000 cm3 C. 0.26 cm3 D. 26 cm3

OpenStudy (skullpatrol):

Same idea, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i don't get this one @SolomonZelman and @skullpatrol

OpenStudy (skullpatrol):